OpenStudy (mariebrittany07):

how do you figure out this problem x-4/3=10

OpenStudy (anonymous):

Just turn it around, 10 + 4/3 = x

OpenStudy (mariebrittany07):

okay how would I solve it or is that all I have to do

OpenStudy (anonymous):

Are you solving for x?

OpenStudy (mariebrittany07):

yes I am solving for x

OpenStudy (mariebrittany07):

I have two more of these to figure out

OpenStudy (anonymous):

So you would need to add 10 + 4/3 to get what x equals. 10 + 4/3 = 34/3 or 11.33333333 So, x=34/3 or 11.3333333

OpenStudy (anonymous):

Haha, didn't you just post another one? where did it go?

OpenStudy (mariebrittany07):

\[\frac{ x-8 }{ 4? }=-10\]

OpenStudy (anonymous):

Why is there a question mark after that 4 in the denominator?

OpenStudy (mariebrittany07):

forgot to take out not suppost to be there

OpenStudy (anonymous):

Ohh, ok. well that was confusing me.. hang on. let me try to work it out.

OpenStudy (mariebrittany07):

okay

OpenStudy (anonymous):

Ok, so I think I'm right.... hahaha. x-8/4=-10 Multiply both sides by the demoninator 4(x-8/4)=(-10)4 The 4's cancel out on the left side and you end up with x-8=-40 Then you add 8 to both sides to get x by itself x=-32

OpenStudy (mariebrittany07):

\[\frac{ x+10 }{ 5 }=8 5(x+10/5)=(8)5 x+10=40 x=30 right

OpenStudy (anonymous):

Is this a new problem? The way you have it written is confusing me.

OpenStudy (mariebrittany07):

i did one by myself just checking to see if i understand yes it is new

OpenStudy (mariebrittany07):

\[\frac{ x+10 }{ 5 }=8\]

OpenStudy (mariebrittany07):

5(x+10/5)=(8)5 x+10=40 x=30 right

OpenStudy (anonymous):

Yay! I think you're right! Good job!

OpenStudy (mariebrittany07):

i have never done algebra before but i need your help on some others

OpenStudy (anonymous):

Ok, just post them & i'll do my best

OpenStudy (mariebrittany07):

this is a new problem 9x+4+2x=37

OpenStudy (mariebrittany07):

9x+2x=11x 11x+4=37 am i right

OpenStudy (anonymous):

So far, yes.. Need help on the rest?

OpenStudy (mariebrittany07):

okay

OpenStudy (anonymous):

It was a question...

OpenStudy (mariebrittany07):

yes

OpenStudy (mariebrittany07):

not sure what to do next

OpenStudy (anonymous):

Ok, so you got to 11x+4=37 Subtract 4 from each side so 11x will by itself.. 11x=33 Then divide both sides by 11 to get x by itself x=3

OpenStudy (mariebrittany07):

okay next question 6x-2-3x=25 6x-3x=3x 3x-2=25

OpenStudy (mariebrittany07):

lost from here

OpenStudy (anonymous):

3x-2=25 Add 2 to both sides 3x=27 Divide both sides by 3 x=9

OpenStudy (mariebrittany07):

okay i understand some

OpenStudy (mariebrittany07):

new problem 4x+5-2x-3=10

OpenStudy (mariebrittany07):

lost have adding and subtraction both

OpenStudy (anonymous):

4x+5-2x-3=10 4x and -2x combine to become 2x 5 and -3 combine to become 2 So... 2x+2=10 Subtract 2 from both sides 2x=8 Divide by 2 on both sides x=4

OpenStudy (mariebrittany07):

you are so smart

OpenStudy (mariebrittany07):

new problem

OpenStudy (mariebrittany07):

4(3x-5)=6 4*3=12x 4*5=20 12x-20=6 am i right

OpenStudy (mariebrittany07):

this is brittany's mom

OpenStudy (anonymous):

How was the original problem stated? & thank you! I am on this site because I need help too but no one can help me..

OpenStudy (mariebrittany07):

dont know math it was not my subject in school

OpenStudy (anonymous):

Hi, Brittany's mom..

OpenStudy (mariebrittany07):

4(3x-5)=6

OpenStudy (mariebrittany07):

hi

OpenStudy (mariebrittany07):

brittany is eatting so i am typing

OpenStudy (mariebrittany07):

you are showing her

OpenStudy (anonymous):

Ok, so start by multiplying out the first part of the problem.. 4(3x-5) 4*3x=12x 4*-5=-20 So now you have this: 12x-20=6 Add 20 to both sides 12x=26 Divide by 12 on both sides x=26/12 then you can simplify that to 13/6 I want you to know.. Although I'm pretty sure I'm right I'm not as confident on this one.

OpenStudy (anonymous):

Ohh, wait.. I am just as confident.. Haha, it's right.. I went back over it.

OpenStudy (mariebrittany07):

this is brittany that was to funny

OpenStudy (mariebrittany07):

-3(-2x-4)=5

OpenStudy (mariebrittany07):

-3*-2=6

OpenStudy (mariebrittany07):

-3*4=-12

OpenStudy (mariebrittany07):

right

OpenStudy (mariebrittany07):

6x- -12=5

OpenStudy (anonymous):

I was about to correct something you did, but your end result that you just posted was right. You know the - - = + right? So 6x+12=5

OpenStudy (mariebrittany07):

so i minus 12 from both sides right giving me a neg 7

OpenStudy (mariebrittany07):

divide -7 by 6

OpenStudy (anonymous):

yes! :)

OpenStudy (mariebrittany07):

x=1.16 right

OpenStudy (anonymous):

Yes and no.. You divided a negative by a positive so your answer has to be negative. So it's -1.167

OpenStudy (mariebrittany07):

okay thank you

OpenStudy (mariebrittany07):

one more

OpenStudy (anonymous):

Ok.. haha

OpenStudy (mariebrittany07):

-4(5x+6)=7 -4*5=-20 -4*6=-24 -20+-24=7

OpenStudy (mariebrittany07):

-20x +-24=7

OpenStudy (anonymous):

So far so good

OpenStudy (mariebrittany07):

this is as far as i know

OpenStudy (anonymous):

Ohh, ok. Then you add 24 to both sides..

OpenStudy (mariebrittany07):

x=31

OpenStudy (anonymous):

No, adding the 24 to both sides didn't effect the -20x. It just got rid of the -24 and made the right side 31.. So at this step you should have: -20x=31 And then you would divide by -20 on both sides

OpenStudy (mariebrittany07):

x=-1.55

OpenStudy (mariebrittany07):

am i right

OpenStudy (anonymous):

Yes! yay

OpenStudy (mariebrittany07):

thank you so much see i know some not all

OpenStudy (mariebrittany07):

you are my lifesaver

OpenStudy (anonymous):

You're welcome! You'll get it! :) Best of luck! I'm getting off line now.. haha I need to try to get back to my statistics.. ughhhhhh.. haha

OpenStudy (mariebrittany07):

okay mom said she wish she could help

OpenStudy (anonymous):

Me too! Haha.. It's okay though, I'll figure it out. Bye (:

OpenStudy (mariebrittany07):

bye